OpenStudy (anonymous):

in 2009, the total population of north Carolina was 2,890,425 people what was the population in scientific notation approximately a 2.9 × 10^5 b 2.9 × 10^6 c 2.9 × 10^7 or d 29 × 10^7 ^= exponent please help

OpenStudy (doc.brown):

\[2,890,425=\]\[2890425\times 1\]\[289042.5\times 10\]\[28904.25\times 10^2\]\[2890.425\times 10^3\]\[289.0425\times 10^4\]\[28.90425\times 10^5\]\[2.890425\times 10^6\]

OpenStudy (doc.brown):

\[10\times 10\times 10\times 10\times 10\times 10 = 10^6\]

OpenStudy (anonymous):

thank you